OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). Answer y = 3x - 5 y = 3x - 2 y = 3x + 1 y = 3x + 3

undeadknight26 (undeadknight26):

i think second one.

OpenStudy (anonymous):

Shanks

OpenStudy (anonymous):

slope(m) = 3 (1,-2) x = 1 and y = -2 y = mx + b ....solving for b -2 = 3(1) + b -2 = 3 + b -2-3 = b -5 = b your line is : y = 3x - 5............first one